Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Methods The units of an item available for sale during the year were as follows: Jan. 1 Inventory10 units at $40 Aug. 7 Purchase 18 units at $41 Dec. 11Purchase 14 units at $42 $400 738 588 $1,726 42 units There are 16 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost using (a) the first-in, first-out (FIFO) method; b) the last-in, first out LIFO method; and c) the weighted average cost method ro nd per unit cost to two dec ma places and your inal answe?to e nearest whole dollar) a. First-in, first-out (FIFO) b. Last-in, first-out (LIFO) C. Weighted average cost
Answer #1

a. FIFO Method

Ending Inventory = 16 Units

As Per FIFO method 2 Units of Aug 7 Purchase and 14 Units of December 11 will be in inventory

Thus Inventory Cost as per FIFO = 2 * $41 + 14 * $42 = $670

b. LIFO Method

Ending Inventory = 16 Units

As Per LIFO method 10 Units of Beginning Inventory and 6 Units of August 7 purchase will be in inventory

Thus Inventory Cost as per LIFO = 10 * $40 + 6 * $41 = $646

c. Weighted Average Method

Weighted Average Cost = Total Cost / No. of Units = $1726 / 42 = $41.10

Ending Inventory = 16 Units

Thus Inventory Cost as per WAM = 16 * $41.095 = $657.60